OpenStudy (scarlettfarra2000):

What is the solution to the system of equations Use the linear combination method 3x-y=4 2x+y=6

OpenStudy (anonymous):

hi ! 3x-y=4 => 3x = 4+y => x= y+4/3 Now use y+4/3 for y in 2x + y = 6 and find y . then use number of y to find x ;) ok ?

OpenStudy (scarlettfarra2000):

thanks

OpenStudy (anonymous):

Did it help ?:p

OpenStudy (scarlettfarra2000):

Yes